OpenStudy (anonymous):

find the slope of the line passing through the points (5,3) and (2,2)?

OpenStudy (anonymous):

\[slope = \frac{y_1-y_2}{x_1-x_2}= \frac{3-2}{5-2}= \frac{1}{3}\]

OpenStudy (anonymous):

Thank you I have one ore please

OpenStudy (anonymous):

oh sure but medal first

OpenStudy (anonymous):

@Smmyjoe2 oh sure but medal first